OpenStudy (ashleyy_smith):

For his phone service, Chris pays a monthly fee of $19, and he pays an additional $0.04 per minute of use. The least he has been charged in a month is $78.16. What are the possible numbers of minutes he has used his phone in a month? Use m for the number of minutes, and solve your inequality for m.

OpenStudy (anonymous):

A

OpenStudy (fortytherapper):

So, we first have to set up an equation. \[y=???? + 19\] Now, for example. Let's say I made 10 dollars an hour and wanted to know how many dollars I would make in 2 hours. It would be 20. Would I multiply, add, subtract, or divide 10 and 2 to get 20?

OpenStudy (ashleyy_smith):

mult @FortyTheRapper

OpenStudy (fortytherapper):

Alright, so we have \[y = 0.04x + 19\] We have an unknown Y(total) and an unknown X(minutes). They actually gave us the total, do you see that?

OpenStudy (ashleyy_smith):

yes

OpenStudy (fortytherapper):

What is it?

OpenStudy (ashleyy_smith):

20?

OpenStudy (fortytherapper):

It's the least he's been charged a month, which is 78.16. With that, we can complete the equation \[0.04m+19 = 78.16\] Now, the last thing we have to do is solve for m, and we can find out the max possible number of minutes he has used in a month

OpenStudy (ashleyy_smith):

oh ok

OpenStudy (fortytherapper):

Let's move the 19 first. What would you do?

OpenStudy (ashleyy_smith):

subtract it by the 78.16? right?

OpenStudy (fortytherapper):

Yep! \[0.04m +19-(19)= 78.16 - 19\] To get us \[0.04m = 59.16\] Now the last step is to get the 0.04 away from the m. Any idea

OpenStudy (ashleyy_smith):

umm divide?

OpenStudy (fortytherapper):

Right \[m = \frac{ 59.16 }{ 0.04}\] Now we just put that in the calculator and there's our answer

OpenStudy (ashleyy_smith):

1,479

OpenStudy (fortytherapper):

Yes, 1,479 minutes. Good job

OpenStudy (ashleyy_smith):

do i turn it into an equation now? cause it says in the end of the problem, use m for the number of minutes, and solve your inequality for m

OpenStudy (fortytherapper):

Nah, those were just instructions on how to do this problem. "What are the possible numbers of minutes he has used his phone in a month" was the question I guess you can say "Between 0 and 1,479 minutes" if you wanted